OpenStudy (anonymous):

Help please Solve each of the following pairs of simultaneous equations for x and y 1) ax+y=c x+by=d

OpenStudy (ranga):

Multiply the second equation by b and then subtract it from the first equation to eliminate x. Solve for y. Then substitute y in the first equation and solve for x.

OpenStudy (anonymous):

@ranga can you give me the work out of it as I don't really understand how it work. Thanks :D

OpenStudy (ranga):

ax + y = c ---- (1) x + by = d ---- (2) multiply (2) by a: ax + aby = ad. Subtract (1): ax + y = c subtract aby - y = ad - c y(ab - 1) = ad - c y = (ad - c) / (ab - 1) Put y in (1) ax + (ad - c) / (ab - 1) = c ax = c - (ad - c) / (ab - 1) = (abc - c - ad + c) / (ab - 1) = (abc - ad) / (ab - 1) ax = a(bc - d) / (ab - 1) divide both sides by a: x = (bc - d) / (ab - 1) x = (bc - d) / (ab - 1) and y = (ad - c) / (ab - 1)

OpenStudy (anonymous):

@ranga thanks :D

OpenStudy (ranga):

You are welcome.